OpenStudy (anonymous):

George and Carmen went on a bicycle trip. They took a bus to their starting point, and then biked the rest. They traveled 300 kilometers in total, and they biked 70 kilometers more than they were bussed. x=kilometers traveled by bike y= kilometers traveled by bus

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (anonymous):

eh wuts ur question

OpenStudy (anonymous):

how many kilometers they traveled by bike

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

bike is x bus is y x + y = 300 x = y + 70 substitute (y + 70) for x in the first equation (y + 70) + y = 300 2y = 230 y = 115 km bussed but you need biked x = 115 + 70 = 185 km biked

OpenStudy (anonymous):

do you agree?

OpenStudy (anonymous):

ya thanks I was having trouble with the formula

OpenStudy (anonymous):

Welcome! c: